OpenStudy (anonymous):

Simplify. 2 times the 4th root of 80

OpenStudy (mathmale):

80 is not a perfect fourth power. What are perfect fourth powers? x x^4 0 0 1 1^4=1 2 2^4=16 3 3^4=81 Split up that 80 so that one factor is a perfect fourth power.

OpenStudy (anonymous):

5 * 16 = 80

OpenStudy (anonymous):

so it would be 4√5??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

times 2?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

so 8 √5

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but it is easy

OpenStudy (mathmale):

Nice work, Hannah!

OpenStudy (anonymous):

Thank you all!